# Properties and dynamics of generalized squeezed states

### v1

@@ III Results
The fact that the small probability that goes up to extremely large values of $m$ consistently returns back close to zero suggests that, because of the large values of $H_{k,k+n}^{(n)}$ at large values of $k$, the large-$k$ part of the Hilbert space acts as a single, strongly-coupled object from the point of view of the small-$k$ part of the Hilbert space. As a result, although the dynamics inside the large-$k$ subspace differs greatly for different truncation conditions, the dynamics between the low and high-$k$ subspaces is truncation independent. Although it is not shown in Fig. 7, we found that the maximum height of the $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ oscillations for $n=4$ also depends on the truncation details. This point is investigated further in Fig. 8(b). By noting that the data follows a straight line on a log-linear plot, we find that $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle_{\rm max}\propto\log[N]$, i.e. weaker dependence that in the case $n=3$. For the case $n=5$, we found that the oscillation height in Fig. 7 is the same for hard and soft cutoff and for $N=5\times 10^{3}$ and $5\times 10^{4}$. We therefore conclude that the $n=5$ curve in Fig. 7 does not contain any finite-size effects. It is quite interesting how the role of finite-size effects changes as we vary $n$: no noticeable effects for $n=1$ and $n=2$, then the clearest effects for $n=3$, milder effects for $n=4$ and back to no noticeable effects for $n=5$.
@@
As mentioned in Sec. II, we do not expect that a realistic near-future experiment will be able to observe a trisqueezed states with an average photon number of hundreds of photons, which is what the above analysis with large values of $N$ predicts. However, since residual terms in a realistic system Hamiltonian could effectively impose a cutoff of only tens or hundreds of photons, it could be possible to observe the oscillatory behaviour of the squeezing dynamics. As the exact behaviour will depend on the details of the actual physical system, we shall not perform an analysis of any specific physical system here.
@@ IV Discussion
An immediate question that comes to mind in this case is whether our results are a mathematical artefact of our numerical simulation approach, especially considering the mathematical complications that were encountered in previous studies on generalized squeezing. We do not believe that this is the case. A strong piece of evidence that our results are not mathematical artefacts is the near-perfect agreement between the simulation results obtained under greatly different conditions. As an example, one can take Fig. 6, where two different simulations produced almost identical results, even though one simulation used 400 basis states with a soft cutoff, while the other used $4\times 10^{4}$ basis states with a hard cutoff. If the results were mathematical artefacts of our truncation, we would expect to see pronounced finite-size effects that are different between these two simulations of vastly different sizes. Another somewhat paradoxical situation is the fact that in Fig. 7 the oscillation height is truncation dependent but the oscillation period is truncation independent. We believe that the latter suggests that the oscillations are not a mathematical artefact.
@@
Figure 10: The difference $\Delta E$ between the two eigenvalues of highest probability in the vacuum state. In Panel (a), we plot $\Delta E$ as a function of squeezing order $n$. Starting from $n=3$, the data clearly follow an exponential scaling law, as illustrated by the solid line, which was obtained by fitting the data to a straight line. The data up to $n=6$ were obtained using $N=n\times 2\times 10^{4}$, after which the two eigenvalues became asymmetric about zero, indicating calculation errors. Therefore, for $n=7$ and $n=8$, we used $N=n\times 2\times 10^{3}$. In Panel (b), we plot $\Delta E$ as a function of truncation size $N$. The inset shows a log-log plot of $\Delta E-3.528$ (which was obtained as a fitting parameter) as a function of $N$. The $x$-axis range is the same as the main plot, and the $y$-axis range is $[6\times 10^{-3},10^{-1}]$. The straight line fit in this log-log plot indicates that $\Delta E$ converges to 3.528 when $N\to\infty$. All data points in this figure were calculated with a hard cutoff.
@@
Figure 11: Probability distribution (as a function of photon number $m$) of the Hamiltonian eigenstate that has the maximum overlap with the vacuum state $\left|0\right\rangle$. The red circles, blue + signs and magenta $\times$ symbols correspond, respectively, to $N=n\times 100$, $n\times 1000$ and $n\times 10^{4}$, all with a hard cutoff. Note that the probability distributions oscillate as functions of $m$, such that there are data points with $P<10^{-4}$ that are not shown in this figure. For $n=1$, the data does not seem to decrease asymptotically with increasing $m$. Furthermore, the probability at each value of $m$ decreases as $1/N$. These two facts show that at least the most important eigenstates are delocalized over the full truncated Hilbert space and can be expected to produce dynamics in which the finite-$r$ state deviates greatly from the initial state $\left|0\right\rangle$. For $n=3$ and $n=4$, the different data sets are almost indistinguishable, which indicates the absence of finite-size effects. For $n=2$, the pattern is different from both those for $n=1$ and $n\geq 3$, which makes it more difficult to make definitive statements based on this figure alone.
@@
To further investigate the difference between the dynamics for different values of $n$, we analyze the eigenstates of the Hamiltonian $\hat{H}_{n}$. Of particular interest is the overlap between the vacuum state $\left|0\right\rangle$, which was used as the initial state in all of our simulations, and the different eigenstates. In Fig. 9, we plot the highest ten eigenstate probabilities in the vacuum state. For $n=1$, the vacuum state is a superposition of a large number of eigenstates, while for $n\geq 3$, the vacuum state is a superposition of mainly two eigenstates. When a quantum system is in a superposition of two Hamiltonian eigenstates, it is natural that the dynamics will exhibit simple oscillations at a frequency given by the difference between the two corresponding eigenvalues. The difference ($\Delta E$) between the two relevant eigenvalues eigenvalue is plotted in Fig. 10. For $n=3$, we find that $\Delta E=3.528$, which gives an oscillation period of $r=2\pi/\Delta E=1.781$, close to the period observed in Figs. 3 and 7. If we plot the eigenvalue difference as a function of $n$ (Fig. 10a), we find that for $n\geq 3$ the data follows an exponential scaling law, which means that the oscillation period decreases rapidly as we increase $n$.
@@
In the same line of investigating Hamiltonian eigenstates, in Fig. 11 we plot the probability distribution of the highest-probability eigenstate as a function of photon number $m$. The contrast between the cases $n=1$ and $n\geq 3$ is clear. In the former case, the probability distribution extends to large numbers of photons, while in the latter case, the probability distribution is strongly localized at small values of $m$. The case $n=2$ lies somewhere in between the extremes of the cases $n=1$ and $n\geq 3$, which makes it difficult to make statements about the case $n=2$ from this eigenstate analysis. Considering the two most relevant eigenstates for $n=3$, they are complex conjugates of each other, and each has $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle=112.6$. If we approximate the initial state as a superposition of only these two states, we expect that $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ will exhibit simple sinusoidal oscillations with amplitude 225. The maximum amplitude for the corresponding curve in Fig. 7 is 320. In other words, the two main eigenstates account for 70% of the oscillation amplitude. Indeed, the irregularity in the oscillation pattern is a clear indication that more than two Hamiltonian eigenstates make substantial contributions. For $n=4$, the main two eigenstates have $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle=10.6$ and hence account for 91% of the oscillation amplitude. For $n=5$, the main two eigenstates have $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle=4.26$ and account for 99% of the oscillation amplitude. We also note here that $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ for the two main eigenstates increases with $N$ following similar scaling laws as those explained in Fig. 8.
@@
One result that we did not show in detail but can be inferred from Fig. 4 is that the probability distribution as a function of photon number $m$ decreases monotonically for small values of $r$ but develops an oscillatory pattern as $r$ is increased. In fact, the oscillations in the long tail of the probability distribution seem to be present even at small values of $r$, but they become noticeable only for larger values of $r$. For example, for $n=3$, the oscillations reach the states with photon number below 30 only for $r>0.2$. This oscillatory pattern in the probability distribution that appears above a certain value of $r$ suggests an interference effect between waves propagating in opposite directions. More specifically, it suggests that the wave of probability amplitudes initially propagates from low to high values of $m$ but is reflected at some point and the interference between the forward and backward propagating waves creates the oscillations. Note that this situation does not occur for displacement ($n=1$) and two-photon squeezing ($n=2$). This situation also raises the question of where the reflection happens. Importantly, if the reflection happens at the end of the photon number ladder in our truncated Hamiltonian, it could be indicative of finite-size effects, and we have to treat the results with caution. If the reflection point is some naturally occurring photon number (which can depend on $n$ but does not depend on $N$), it would be an indication that this reflection is a physical result as opposed to a mathematical artefact. The fact that the probability pattern depends only weakly on $N$, which covers three orders of magnitude in our simulations, as shown in Fig. 6, leads us to believe that the reflection is a physical phenomenon rather than being a finite-size effect.
@@
The function $x(t)$ for $n=1$ exhibits the correct quadratic growth of the position (which represents the number of photons) with time, while the functions for $n\geq 2$ do not accurately describe the quantum dynamics. It is also interesting to note that in the cases $n=1$ and $n=2$ the position grows indefinitely up to $t\to\infty$, while for $n\geq 3$, it diverges at a finite value of $t$, indicating a qualitative difference between the cases $n\leq 2$ and $n\geq 3$.
@@
Figure 12: Q and Wigner functions for quantum states that are constructed by taking the amplitudes of the states $\left|0\right\rangle$, $\left|2\right\rangle$, $\left|4\right\rangle$ in the state $\hat{U}_{2}(r)\left|0\right\rangle$ and assigning these amplitudes to the states $\left|0\right\rangle$, $\left|3\right\rangle$, $\left|6\right\rangle$ In other words, these are artificially constructed trisqueezed states. The squeezing parameter is given by $r=0.1$ (top row), 0.5 (middle row) and 1.0 (bottom row).
@@
We finally note that one can generate squeezed states whose Q and Wigner functions keep getting elongated indefinitely, even for $n\geq 3$, by using different Hamiltonians. For example, we can take inspiration from the dynamics of two-photon squeezing and design a Hamiltonian whose matrix elements vanish except for $H_{3k,3(k+1)}=\sqrt{(2k+1)(2k+2)}$ for all integer values of $k$. If we now take this Hamiltonian and the initial state $\left|0\right\rangle$, we find that the probability amplitudes of the states $\left|0\right\rangle,\left|3\right\rangle,\left|6\right\rangle,...$ grow in exactly the same way that the amplitudes of the states $\left|0\right\rangle,\left|2\right\rangle,\left|4\right\rangle,...$ grow with two-photon squeezing. As a result, the three arms of the trisqueezed state keep growing indefinitely. An example of such states is shown in Fig. 12. Another way to look at the possibility of designing indefinitely growing generalized squeezed states is as a case of reverse engineering: if we know the quantum state evolution that we want to induce, we can differentiate the state and substitute it in the Schrödinger equation to obtain the necessary Hamiltonian.

### v2

@@ III Results
To investigate this point further, in Fig. 8 we plot the maximum value of $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ reached in the first oscillation as a function of matrix size $N$. For $n=3$, by fitting the data in Fig. 8(a) to a straight line, we find that to a good approximation $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle_{\rm max}\propto N^{0.562}$ for the case of a hard cutoff and $N^{0.610}$ for the case of a soft cutoff. These functions diverge as $N\to\infty$. However, it should be noted that the divergence is sub-linear. Most of the probability remains at small photon numbers, and the probability that goes to the maximum photon number decreases with increasing $N$. Still, the small probability in the tail of the distribution is sufficient to raise the average photon number to infinity in the limit $N\to\infty$. Although it is not shown in Fig. 7, we found that the maximum height of the $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ oscillations for $n=4$ also depends on the truncation details. This point is investigated further in Fig. 8(b). By noting that the data follows a straight line on a log-linear plot, we find that $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle_{\rm max}\propto\log[N]$, i.e. weaker dependence that in the case $n=3$. For the case $n=5$, we found that the oscillation height in Fig. 7 is the same for hard and soft cutoffs and for $N=5\times 10^{3}$ and $5\times 10^{4}$. We therefore conclude that the $n=5$ curve in Fig. 7 does not contain any finite-size effects. It is quite interesting how the role of finite-size effects changes as we vary $n$: no noticeable effects for $n=1$ and $n=2$, then the clearest effects for $n=3$, milder effects for $n=4$ and back to no noticeable effects for $n=5$.
@@
As mentioned in Sec. II, we do not expect that a realistic near-future experiment will be able to observe a trisqueezed states with an average photon number of hundreds of photons, which is what the above analysis with large values of $N$ predicts. For example, Ref. Eriksson reached $r=0.13$ for $n=3$, while Ref. Bazavan reached $r=0.19$ for $n=3$ and $r=0.054$ for $n=4$. For $n=2$, the record squeezing parameter for light is $r=1.7$ Vahlbruch . These numbers cast doubt on the possibility of observing any effects that involve large photon numbers. However, since residual terms in a realistic system Hamiltonian could effectively impose a cutoff of only tens or hundreds of photons, it should be possible to observe the oscillatory behaviour of the squeezing dynamics without the need to reach high photon numbers. As the exact behaviour will depend on the details of the actual physical system, we shall not perform an analysis of any specific physical system here.
@@ IV Discussion
In this section we discuss a few additional aspects of the results presented in Section III.
IV.1 Mathematical aspects of truncated simulations
@@
An immediate question that comes to mind in this case is whether our results are a mathematical artefact of our numerical simulation approach, especially considering the mathematical complications that were encountered in previous studies on generalized squeezing. A strong piece of evidence that our results are not mathematical artefacts is the near-perfect agreement between the simulation results obtained under greatly different conditions. As an example, one can take Fig. 6, where two different simulations produced almost identical results, even though one simulation used 400 basis states with a soft cutoff, while the other used $4\times 10^{4}$ basis states with a hard cutoff. If the results were mathematical artefacts of our truncation, we would expect to see pronounced finite-size effects that are different between these two simulations of vastly different sizes. Another somewhat paradoxical situation is the fact that in Fig. 7 the oscillation height is truncation dependent but the oscillation period is truncation independent. The latter suggests that the oscillations are not a mathematical artefact.
In this context, we also note another remarkable aspect of the fact that the average photon number grows indefinitely with increasing $N$. Normally, when a simulated quantity reaches the edge of the simulation space and comes back, it signals that the simulation results cannot be trusted after the reflection off the artificially imposed boundary of the simulation space. For example, we encounter this situation in the case $n=2$ if we keep increasing $r$. In that case, however, the $r$ value at which the state comes back close to the initial state also grows indefinitely with increasing $N$. In the case of the $n=3$ oscillations, the system consistently goes back to the vacuum state at $r=1.75$. One intuitive way to think about the divergence and subsequent decrease of the average photon number is as follows: because of the large values of $H_{k,k+n}^{(n)}$ at large values of $k$, the large-$k$ part of the Hilbert space can be thought of as a single, strongly-coupled subspace as seen from the small-$k$ subspace. As a result, although the dynamics inside the large-$k$ subspace differs greatly for different truncation conditions, the dynamics between the low and high-$k$ subspaces is almost truncation independent.
IV.2 Eigenvalues and eigenstates of squeezing generator
@@
Figure 10: The difference $\Delta E$ between the two eigenvalues that have the highest probabilities in the vacuum state. In Panel (a), we plot $\Delta E$ as a function of squeezing order $n$. Starting from $n=3$, the data clearly follow an exponential scaling law, as illustrated by the solid line, which was obtained by fitting the data to a straight line. The data up to $n=6$ were obtained using $N=n\times 2\times 10^{4}$, after which the two eigenvalues became asymmetric about zero, indicating calculation errors. Therefore, for $n=7$ and $n=8$, we used $N=n\times 2\times 10^{3}$. In Panel (b), we plot $\Delta E$ as a function of truncation size $N$ for $n=3$. The inset shows a log-log plot of $\Delta E-3.528$ (which was obtained as a fitting parameter) as a function of $N$. The $x$-axis range is the same as the main plot, and the $y$-axis range is $[6\times 10^{-3},10^{-1}]$. The straight line fit in this log-log plot indicates that $\Delta E$ converges to 3.528 when $N\to\infty$. All data points in this figure were calculated with a hard cutoff.
@@
Figure 11: Probability distribution (as a function of photon number $k$) of the Hamiltonian eigenstate that has the maximum overlap with the vacuum state $\left|0\right\rangle$. The red circles, blue + signs and magenta $\times$ symbols correspond, respectively, to $N=n\times 100$, $n\times 1000$ and $n\times 10^{4}$, all with a hard cutoff. Note that the probability distributions oscillate as functions of $k$, such that there are data points with $P<10^{-4}$ that are not shown in this figure. For $n=1$, the data does not seem to decrease asymptotically with increasing $k$. Furthermore, the probabilities decrease significantly with increasing $N$. These two facts show that the most important eigenstates are delocalized over the full truncated Hilbert space. For $n=3$ and $n=4$, the probabilities decrease quickly as functions of $k$ and exhibit only weak dependence on $N$. For $n=2$, the pattern is different from both those for $n=1$ and $n\geq 3$, which makes it more difficult to make definitive statements based on this figure alone.
@@
To further investigate the difference between the dynamics for different values of $n$, we analyze the eigenstates of the Hamiltonian $\hat{H}_{n}$. Of particular interest is the overlap between the vacuum state $\left|0\right\rangle$, which was used as the initial state in all of our simulations, and the different eigenstates. In Fig. 9, we plot the highest ten eigenstate probabilities in the vacuum state. For $n=1$, the vacuum state is a superposition of a large number of eigenstates, while for $n\geq 3$, the vacuum state is a superposition of mainly two eigenstates. When a quantum system is in a superposition of two Hamiltonian eigenstates, it is natural that the dynamics will exhibit simple oscillations at a frequency given by the difference between the two corresponding eigenvalues. The difference ($\Delta E$) between the two most relevant eigenvalues is plotted in Fig. 10. For $n=3$ (analyzed in detail in Fig. 10b), we find that $\Delta E=3.528$, which gives an oscillation period of $r=2\pi/\Delta E=1.781$, close to the period observed in Figs. 3 and 7. From the plot of the eigenvalue difference $\Delta E$ as a function of $n$ (Fig. 10a), we find that for $n\geq 3$ the data follows an exponential scaling law, which means that the oscillation period decreases rapidly as we increase $n$.
@@
In the same line of investigating Hamiltonian eigenstates, in Fig. 11 we plot the probability distribution of the highest-probability eigenstate as a function of photon number $k$. The contrast between the cases $n=1$ and $n\geq 3$ is clear. In the former case, the probability distribution extends to large values of $k$, while in the latter case, the probability distribution is strongly localized at small values of $k$. The case $n=2$ lies somewhere in between the extremes of the cases $n=1$ and $n\geq 3$, which makes it difficult to make statements about the case $n=2$ from this eigenstate analysis. Considering the two most relevant eigenstates for $n=3$, they are complex conjugates of each other, and each has $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle=112.6$. If we approximate the initial state as a superposition of only these two states, we expect that $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ will exhibit simple sinusoidal oscillations with amplitude 225. The maximum amplitude for the corresponding curve in Fig. 7 is 320. In other words, the two main eigenstates account for 70% of the oscillation amplitude. Indeed, the irregularity in the oscillation pattern is a clear indication that more than two Hamiltonian eigenstates make substantial contributions. For $n=4$, the main two eigenstates have $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle=10.6$ and hence account for 91% of the oscillation amplitude. For $n=5$, the main two eigenstates have $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle=4.26$ and account for 99% of the oscillation amplitude. We also note here that $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle$ for the two main eigenstates increases with $N$ following similar scaling laws as those explained in Fig. 8.
IV.3 Analogy with wave propagation
@@
One result that we did not show in detail but can be inferred from Fig. 4 is that the probability distribution as a function of photon number $k$ decreases monotonically for small values of $r$ but develops an oscillatory pattern as $r$ is increased. In fact, the oscillations in the long tail of the probability distribution seem to be present even at small values of $r$, but they become noticeable only for larger values of $r$. For example, for $n=3$, the oscillations reach the states with photon number below 30 only for $r>0.2$. This oscillatory pattern in the probability distribution that appears above a certain value of $r$ suggests an interference effect between waves propagating in opposite directions. More specifically, it suggests that the wave of probability amplitudes initially propagates from low to high values of $k$ but is reflected at some point and the interference between the forward and backward propagating waves creates the oscillations. Note that this situation does not occur for displacement ($n=1$) and two-photon squeezing ($n=2$). This situation also raises the question of where the reflection happens. Importantly, if the reflection happens at the end of the photon number ladder in our truncated Hamiltonian, it could be indicative of finite-size effects, and we have to treat the results with caution. If the reflection point is some naturally occurring photon number (which can depend on $n$ but does not depend on $N$), it would be an indication that this reflection is a physical result as opposed to a mathematical artefact. The fact that the probability pattern depends only weakly on $N$, which covers three orders of magnitude in our simulations, as shown in Fig. 6, leads us to believe that the reflection is a physical phenomenon rather than being a finite-size effect.
@@ IV.4 Classical analogy
The function $x(t)$ for $n=1$ exhibits the correct quadratic growth of the position (which represents the number of photons) with time, while the functions for $n\geq 2$ do not accurately describe the quantum dynamics. It is nevertheless interesting to note that in the cases $n=1$ and $n=2$ the position grows indefinitely up to $t\to\infty$, while for $n\geq 3$, it diverges at a finite value of $t$ (assuming that $x(0)\neq 0$ for $n\geq 2$), indicating a qualitative difference between the cases $n\leq 2$ and $n\geq 3$.
IV.5 Squeezing using the operator $\hat{p}^{n}$
Figure 12: Q and Wigner functions for the squeezed vacuum state obtained using the squeezing operator in Eq. (11) with $n=3$, also known as cubic phase state, for $r=0.1$, 0.3 and 0.5. Unlike the trisqueezed state discussed in Sec. III, the cubic phase state does not have three-fold symmetry, and it does not undergo periodic dynamics.
It is interesting to consider the alternative effective Hamiltonian Gottesman
$\hat{H}_{n}=\left(\frac{\hat{a}^{\dagger}-\hat{a}}{i}\right)^{n}.$ (11)
Since this operator can be expressed as $\hat{p}^{n}$, where $\hat{p}$ is a rescaled momentum operator, it is possible to perform analytical derivations in this case, e.g. using wave functions in momentum space. One can, with some algebra, show that the average photon number in the squeezed vacuum state is given by
$\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle=(2n-3)!!n^{2}r^{2},$ (12)
where $m!!=m(m-2)\cdots 1$. In other words, the squeezed state keeps growing and does not undergo periodic dynamics. It should be noted, however, that these squeezed states are qualitatively different from those generated by the squeezing operator in Eq. (6). The operator in Eq. (11) contains terms that create or annihilate different numbers of photons. For example, for $n=3$, there are single-photon terms. As a result, the squeezed states do not have three-fold rotation symmetry anymore, as can be seen in Fig. 12. Similar states were reported in Ref. Eriksson . It is interesting to note here that the average photon number in Eq. (12) always grows quadratically in $r$, while $n$ determines the prefactor in the formula.
It should also be noted that the operator in Eq. (6) arises more naturally in experiment than the operator in Eq. (11). The effective Hamiltonian that generates the squeezing is typically derived from the Hamiltonian
$\hat{H}=\hbar\omega\hat{a}^{\dagger}\hat{a}+\Omega\left(\frac{\hat{a}^{\dagger}-\hat{a}}{i}\right)^{n}\sin n\omega t.$ (13)
Upon making the rotating-wave approximation, the only terms that are time independent in the rotating frame will be ones containing $(\hat{a}^{\dagger})^{n}$ and $\hat{a}^{n}$, leading to the effective Hamiltonian in Eq. (6).
IV.6 Avoiding the oscillations with generalized squeezing
@@
Figure 13: Q and Wigner functions for quantum states that are constructed by taking the amplitudes of the states $\left|0\right\rangle$, $\left|2\right\rangle$, $\left|4\right\rangle$ in the state $\hat{U}_{2}(r)\left|0\right\rangle$ and assigning these amplitudes to the states $\left|0\right\rangle$, $\left|3\right\rangle$, $\left|6\right\rangle$ In other words, these are artificially constructed trisqueezed states. The squeezing parameters used to generate these plots are given by $r=0.1$ (top row), 0.5 (middle row) and 1.0 (bottom row).
@@
We finally note that one can generate squeezed states whose Q and Wigner functions keep getting elongated indefinitely, even for $n\geq 3$, by using different Hamiltonians. For example, we can take inspiration from the dynamics of two-photon squeezing and design a Hamiltonian whose matrix elements vanish except for $H_{3k,3(k+1)}=\sqrt{(2k+1)(2k+2)}$ for all integer values of $k$. If we now take this Hamiltonian and the initial state $\left|0\right\rangle$, we find that the probability amplitudes of the states $\left|0\right\rangle,\left|3\right\rangle,\left|6\right\rangle,...$ grow in exactly the same way that the amplitudes of the states $\left|0\right\rangle,\left|2\right\rangle,\left|4\right\rangle,...$ grow with two-photon squeezing. As a result, the three arms of the trisqueezed state keep growing indefinitely. An example of such states is shown in Fig. 13.
Another way to look at the possibility of designing indefinitely growing generalized squeezed states is as a case of reverse engineering: if we know the quantum state evolution that we want to induce, we can differentiate the state and substitute it in the Schrödinger equation to obtain the necessary Hamiltonian. Our results demonstrate that the intuitive, straightforward generalization of displacement and squeezing protocols cannot be used to obtained arbitrarily squeezed states with $n\geq 3$. New protocols must be developed for this purpose. For near-future applications, e.g. for photon numbers below 100, optimal control theory (OCT) can be a powerful method to design control protocols that achieve the desired goal. For example, OCT was used to generate arbitrary few-photon states in a superconducting resonator controlled via a qubit Heeres .